# Multi-Source Music Generation with Latent Diffusion

### v1

@@ Abstract
Most music generation models directly generate a single music mixture. To allow for more flexible and controllable generation, the Multi-Source Diffusion Model (MSDM) has been proposed to model music as a mixture of multiple instrumental sources (e.g. piano, drums, bass, and guitar). Its goal is to use one single diffusion model to generate consistent music sources, which are further mixed to form the music. Despite its capabilities, MSDM is unable to generate songs with rich melodies and often generates empty sounds. Also, its waveform diffusion introduces significant Gaussian noise artifacts, which compromises audio quality. In response, we introduce a multi-source latent diffusion model (MSLDM) that employs Variational Autoencoders (VAEs) to encode each instrumental source into a distinct latent representation. By training a VAE on all music sources, we efficiently capture each source’s unique characteristics in a source latent that our diffusion model models jointly. This approach significantly enhances the total and partial generation of music by leveraging the VAE’s latent compression and noise-robustness. The compressed source latent also facilitates more efficient generation. Subjective listening tests and Fréchet Audio Distance (FAD) scores confirm that our model outperforms MSDM, showcasing its practical and enhanced applicability in music generation systems. We also emphasize that modeling sources is more effective than direct music mixture modeling. Codes and models are available at https://github.com/XZWY/MSLDM. Demos are available at https://xzwy.github.io/MSLDMDemo/.
@@ I Introduction
Generative models show impressive performance not only in language and image modeling [1, 2, 3], but also show great results in music generation. Music generation models usually fall into two categories: 1) Auto-regressive models and 2) Diffusion models.
@@
For auto-regressive models, WaveNet [4] directly models scalar-quantized waveform samples, which allows for generating small musical fragments. However, due to the sample-level auto-regression, WaveNet has low sampling efficiency. One way to improve efficiency is to encode waveform samples to a discrete latent representation (tokens) with a much lower time resolution. These tokenizers [5, 6, 7] are usually variations of VQ-VAE [8], but are usually trained with perceptual adversarial loss [6, 7, 9, 10]. Then the auto-regressive transformer models the sequence of tokens achieving higher efficiency. Among these models, JukeBox [5] allows music generation conditioned on lyrics. More recently, text-to-music generation has shown significant progress based on this framework [11, 12].
@@
On the other hand, diffusion-based music generation models also have great potential, where the diffusion model is learned on some intermediate representation. Noise2Music [13] uses diffusion models to generate an intermediate representation, either downsampled waveforms, or MEL-Spectrogram features, and then decodes the intermediate representation to music waveform by a cascader or vocoder. Then, because of the success of latent diffusion in image generation [14], music generation also starts to use latent diffusion. DiffSound [15] first trains a spectrogram VQ-VAE tokenizer as the intermediate representation, and then uses a discrete diffusion model to model the token sequence. [16, 17, 18, 19] are using spectrogram-domain (variational) autoencoder’s continuous latent as an intermediate representation for diffusion, while StableAudio [20] uses waveform-domain VAE’s latent as the diffusion target. Moûsai [21] proposes a spectrogram encoder learned by diffusion magnitude autoencoding (DMAE) and then trains another diffusion model on the encoder’s latent. Further, with a waveform-domain VAE, StableAudio2 [22] achieves full-song generation by modeling the VAE latent with diffusion.
@@
Although rapid progress has been made in music generation, most models directly model the whole music piece, which is a mixture of individual sources. However, the individual sources cannot be disentangled from the mixture. Ideally, a music generation method should be able to generate music sources that together form a piece of music, similar to a human’s music composition process. This will allow the music to be more interpretable and controllable (e.g., the piano can be made louder than the drums). To solve this problem, one class of methods directly models the musical notes or midi representations in a multi-track manner [23, 24]. However, the generated notes or midi sequence need to be later decoded to a single waveform using synthesizers. The other type of research directly learns to model several music tracks directly. StemGen [25] uses a masked language model on Encodec tokens to generate any single instrument source given a music context. [26] uses the latent diffusion model to generate bass companions conditioned on mixtures, while SingSong [27] generates background companions given the vocal source. Most recently, MSDM [28] has proposed to simultaneously model four instrument sources (piano, drums, bass, guitar) with a single waveform-domain diffusion model, and GMSDI [29] has generalized MSDM by training on text-conditioned diffusion models allowing adaption to any music dataset using text descriptions.
@@ II Models
Fig. 1 shows the training and inference pipeline of our model. Just like any latent diffusion model, our model involves two blocks: 1) SourceVAE which is trained like a VAE but with adversarial loss for perceptual compression. 2) A diffusion model simultaneously models all the sources’ latents concatenated together as one latent. Inference includes two sub-tasks as well: 1) Total generation allows unconditional generation of all instrumental sources at the same time and 2) Partial generation allows the generation of companion sources given any combinations of instrumental sources (e.g., generate bass and guitar to accompany given piano and drums).
@@ II-A SourceVAE
The SourceVAE aims to compress waveform-domain instrumental sources into a compact latent space, while still ensuring perceptually indistinguishable reconstruction. This is usually achieved by adversarial training with carefully designed discriminators. We borrow this training framework and model architecture from the DAC [10] neural audio codec. DAC is a state-of-the-art waveform-domain neural audio codec trained with both reconstruction and adversarial losses. It can encode, quantize, and decode the audio with superior quality. However, we want our latent to be noise-robust and continuous, so we remove the vector quantization module, constrain the intermediate latent size, and add a small KL-divergence loss term as used in vanilla VAEs [30]. The KL-divergence loss is to ensure a noise-robust latent space. For the encoder and decoder, we use DAC’s default 24kHz model but with a new latent size of $C=80$. Given an instrumental source $s\in\mathbb{R}^{N}$ with $N$ samples, the encoder encodes the waveform to a posterior $\Psi_{enc}(s)=\mathcal{N}(\cdot|\mu_{z}(s),\Sigma_{z}(s))$, where $\mu_{z}(s)\in\mathbb{R}^{C\times\frac{N}{D}}$ is the posterior mean of the latent and $\Sigma_{s}$ is the corresponding posterior covariance. $D$ is the time-domain downsampling factor of the encoder, which is $320$ in DAC. Then any $z\sim\mathcal{N}(\cdot|\mu_{z}(s),\Sigma_{z}(s))$, processed by the decoder $\psi_{\text{dec}}$, should reconstruct $s$ with good quality. To extract latent features, we just sample $z_{s}=\mu_{z}(s)$. During training, the loss is as shown below:
@@
where $L_{\text{Mel}}$, $L_{\text{feature}}$, $L_{\text{adversarial}}$ are Mel-reconstruction loss, feature matching loss, and adversarial loss respectively, as in the DAC training framework. Also $\lambda_{1}=15$, $\lambda_{2}=2$, $\lambda_{3}=1$ according to DAC. $L_{\text{KL}}$ is the KL-Divergence between the VAE encoded posterior and $\mathcal{N}(0,I)$ and we set $\lambda=10$. The implementation details are available in our source code.
@@ II-C1 Total Generation
The total generation inference is straightforward. Starting from randomly sampled white noise $Z(T)\sim\mathcal{N}(\cdot|0,\sigma_{max}^{2}I)$, the diffusion sampling gradually transforms $Z(T)$ to $Z(0)$. Then instrumental source latent $z_{1},z_{2},...,z_{K}$ are extracted from $Z(0)$, and are further decoded independently by the SourceVAE decoder to get the generated source waveforms $\{s_{i}\in\mathbb{R}^{N}|s_{i}=\psi_{dec}(z_{i}),i\in[1,2,...,K]\}$. These generated sources could then be added to form a mixture of music pieces $x=\sum_{k=1}^{K}s_{i}$.
@@ II-C2 Partial Generation
Partial generation is the task of generating complementary sources given some existing ones to condition on. Assume a subset of instruments are given by the indices $I\subset\{1,2,...K\}$, and the corresponding given sources are denoted by $S_{I}=\{s_{i}\}_{i\in I}$. Then the complementary sources to generate are indexed by $\bar{I}=\{1,...,K\}\backslash I$, which means the source to generate are $S_{\bar{I}}=\{s_{i}\}_{i\in\bar{I}}$. Since our diffusion model works in the latent domain, we first use SourceVAE to encode each source in $S_{I}$ to latent $Z_{I}=\{z_{i}|z_{i}=\mu_{z}(s_{i}),i\in I\}$. The the task is to generate $Z_{\bar{I}}$ conditioned on $Z_{I}$, so we need the conditional score $\nabla_{{Z}_{\bar{I}}(t)}\log p(Z_{\bar{I}}(t)|Z_{I}(t))$ for sampling. Following diffusion-based imputation [31] and MSDM [28], we could estimate the condition score by:
@@
$d{Z_{\bar{I}}}(t)=\sigma(t)S^{\theta}([Z_{\bar{I}}(t),\hat{Z}_{I}(t)],\sigma^{2}(t))dt$ (6)
With ${Z_{\bar{I}}}(0)$ sampled, then similar to total generation, the partially generated sources $S_{\bar{I}}$ can be decoded by the SourceVAE decoder: $S_{\bar{I}}=\{s_{i}\in\mathbb{R}^{N}|s_{i}=\psi_{dec}(z_{i}),z_{i}\in Z_{\bar{I}}\}$. Then all the conditional sources and partially generated sources are added together to form the final music piece $x$.
@@ III-A Dataset
We use the same dataset as MSDM [28], which is a subset of the slakh2100 music dataset [33]. slakh2100 is a MIDI synthesized music dataset with 145 hours of music, containing both the mixed music and the individual tracks labeled with instrument class. Same as MSDM, we use the $K=4$ main tracks which are piano, drums, bass, and guitar for multi-source modeling. For a fair comparison with MSDM, we use the same sampling rate of 22050Hz.
@@ III-B SourceVAE
The SourceVAE mentioned in Sec. II-A is a 1D-CNN-based encoder-decoder architecture coupled with some DAC loss and KL-divergence loss. The encoder and decoder all follow the final setup of the DAC architecture for 24kHz. The intermediate latent dimension for SourceVAE is set to be $C=80$ and the encoder has a downsampling rate of $D=320$ for the temporal dimension. For training, we use a batch size of 28 and train on one-second-long single-instrumental segment for 10k steps. All other SourceVAE training configurations are the same as the DAC paper with code in https://github.com/descriptinc/descript-audio-codec.
@@ III-C Latent Diffusion and Unet Architecture
To accommodate the 1D-Unet architecture used in MSDM, we concatenate the K source latent on the same channel dimension to form a new channel dimension, so the new channel dimension becomes $KC$, and the input to the 1D-Unet is $Z^{\prime}(t)\in\mathbb{R}^{KC\times\frac{N}{D}}$, which is a reshaped version of $Z(t)$. $KC$ is treated as the channel dimension and $\frac{N}{D}$ is treated as the temporal dimension for the Unet to process. In our setup, $K=4,C=80,D=320$. We train our diffusion model on segments with $N=327672$ samples (about 15 seconds). For the 1D-Unet architecture, similar to MSDM, we adapt the architecture used in Moûsai [21] but with some modifications. We set the input channel dimension to be $KC=320$ and then we experiment on two different configurations. We call one model MSLDM and one larger model MSLDM-Large. The MSLDM contains 6 nested U-Net blocks with increasing channels [1024, 2048, 4096, 4096, 4096, 4096]. The downsampling factor for all the blocks is [1,1,2,1,1,2]. The self-attention blocks are used for all the blocks except the first one. 12 Attention heads are used, and each head is 64 dimensional. For MSDLM-Large, the Unet contains 8 layers where the corresponding output channel dimensions are [1024, 2048, 4096, 4096, 4096, 4096, 4096, 4096]. The corresponding downsampling factors are [1,1,2,1,1,2,2,2]. All blocks contain self-attention blocks except the first block and the attention block contains 12 attention heads where each is 128 dimensional. The Unet and diffusion code setup is adapted from audio-diffusion-pytorch. Similar to MSDM, We train the diffusion with a batch size 16 and a learning rate of 2e-5 for 400k steps.
All the model training is done on a single RTX A6000 GPU with 48GB of VRAM. Further details can be found in our code.
@@ IV Evaluation Metrics and Results
We evaluate the two tasks: 1) Total generation and 2) Partial generating as mentioned in Sec. II-C. For both tasks, we evaluate the FAD score [34] and human subjective score with a listening test. We also design a few baseline models for fair comparison.
@@ IV-A Baseline Models
The first baseline is the MSDM [28] model, which we directly use the trained checkpoints. To show the effectiveness of modeling sources instead of mixtures, we design another baseline called MixLDM, where the latent diffusion model directly models the generation of the latent of music mixture, instead of instrumental sources. For training this model, we first train a MixtureVAE which is the same as SourceVAE except that it’s trained on Mixture music and the latent size $C=320$, so that the diffusion model’s input is of the same shape with MSLDM. This is a more common practice in diffusion-based audio/music generation, where the model directly models the music mixture. To claim that our model generates sources that are consistent or in harmony, we design one baseline called ISLDM (Independent Source Latent Diffusion Model), where we train four independent diffusion models on four instruments’ latents respectively, so each model can generate one single instrument. For each SSLDM model, the input channel size becomes $C=80$, and all the other training is the same as MSLDM. Then to generate a mixture, all four models need to generate one instrumental source, and then these sources could be added together to form the mixture. However, all the sources are generated independently which means the generated mixture would not be in harmony. All models’ parameters and inference time to generate one 12-second mixture are listed in Table I.
@@
MSLDM 364 5.44
MSLDM-Large 1654 7.44
@@
TABLE II: Sub-FAD for Partial Generation. The sub-FAD (lower is better) is reported for any source combinations (B: Bass, D: Drums, G: Guitar, P: Piano), where BD means conditioned on a mixture of piano and guitar, the task is to generate Bass and Drums.
@@
IV-B FAD
We use the Fréchet Audio Distance (FAD) [34] as the objective metric to evaluate total generation and partial generation. We use 200 chunks of music segment for the FAD calculation, where each segment is about 12 seconds, as in MSDM [28]. We randomly sample 200 12-second long music segments from the slakh2100 test set as the reference set.

### v3

@@ Abstract
Most music generation models directly generate a single music mixture. To allow for more flexible and controllable generation, the Multi-Source Diffusion Model (MSDM) has been proposed to model music as a mixture of multiple instrumental sources (e.g. piano, drums, bass, and guitar). Its goal is to use one single diffusion model to generate mutually-coherent music sources, that are then mixed to form the music. Despite its capabilities, MSDM is unable to generate music with rich melodies and often generates empty sounds. Its waveform diffusion approach also introduces significant Gaussian noise artifacts that compromise audio quality. In response, we introduce a Multi-Source Latent Diffusion Model (MSLDM) that employs Variational Autoencoders (VAEs) to encode each instrumental source into a distinct latent representation. By training a VAE on all music sources, we efficiently capture each source’s unique characteristics in a “source latent”. The source latents are concatenated and our diffusion model learns this joint latent space. This approach significantly enhances the total and partial generation of music by leveraging the VAE’s latent compression and noise-robustness. The compressed source latent also facilitates more efficient generation. Subjective listening tests and Fréchet Audio Distance (FAD) scores confirm that our model outperforms MSDM, showcasing its practical and enhanced applicability in music generation systems. We also emphasize that modeling sources is more effective than direct music mixture modeling. Codes and models are available at https://github.com/XZWY/MSLDM. Demos are available at https://xzwy.github.io/MSLDMDemo/.
@@ I Introduction
Generative models show impressive performance not only in language and image modeling [1, 2, 3], but also show promising results in music generation. Music generation models usually fall into two categories: 1) Auto-regressive models and 2) Diffusion models.
@@
For auto-regressive models, WaveNet [4] directly models scalar-quantized waveform samples, which allows for generating small musical fragments. However, due to the sample-level auto-regression, WaveNet has low sampling efficiency. One way to improve efficiency is to encode waveform samples to a discrete latent representation (tokens) with a much lower time resolution. These tokenizers [5, 6, 7] are usually variations of VQ-VAE [8], but are usually trained with perceptual adversarial loss [6, 7, 9, 10]. Then, the auto-regressive transformer models the sequence of tokens achieving higher efficiency. Among these models, JukeBox [5] allows music generation conditioned on lyrics. More recently, text-to-music generation has shown significant progress based on this framework [11, 12].
@@
On the other hand, diffusion-based music generation models also hold great potential, where the diffusion model is learned on some intermediate representation. Noise2Music [13] uses diffusion models to generate an intermediate representation, either downsampled waveforms, or Mel-Spectrogram features, and then decodes the intermediate representation to music waveform by a cascader or vocoder. Then, because of the success of latent diffusion in image generation [14], music generation also follows this path. DiffSound [15] first trains a spectrogram VQ-VAE tokenizer as the intermediate representation, and then uses a discrete diffusion model to model the token sequence. [16, 17, 18, 19] are using spectrogram-domain (variational) autoencoder’s continuous latent as an intermediate representation for diffusion, while [20, 21] uses waveform-domain VAE’s latent as the diffusion target. Moûsai [22] proposes a spectrogram encoder learned by diffusion magnitude autoencoding (DMAE) and then trains another diffusion model on the encoder’s latent. Further, with a waveform-domain VAE, StableAudio2 [23] achieves full-song generation by modeling the VAE latent with diffusion.
@@
Although rapid progress has been made in music generation, most models directly generate the whole music piece, which is a mixture of individual sources. However, the individual sources cannot be disentangled from the mixture. Ideally, a music generation method should be able to generate the individual music sources that together form a piece of music, similar to a human’s music composition process. This will allow the music to be more interpretable and controllable (e.g., the piano can be made louder than the drums). To solve this problem, one class of methods directly models the musical notes or midi representations in a multi-track manner [24, 25]. However, the generated notes or midi sequence need to be later decoded to a single waveform using synthesizers. The other type of research directly learns to model several music tracks directly. StemGen [26] uses a masked language model on Encodec tokens to generate any single instrument source given a music context. [27] uses the latent diffusion model to generate bass companions conditioned on mixtures, while SingSong [28] generates background companions given the vocal source. Most recently, MSDM [29] has been proposed to simultaneously model four instrument sources (piano, drums, bass, guitar) with a single waveform-domain diffusion model, and GMSDI [30] has generalized MSDM by training on text-conditioned diffusion models allowing adaption to any music dataset using text descriptions. The closest work to ours is multi-track MusicLDM, [31, 32], which is simultaneous to this paper, some implementation differences exist.
@@ II Models
Fig. 1 shows the training and inference pipeline of our model. Just like any latent diffusion model, our model involves two blocks: (1) SourceVAE, which is trained like a VAE but with an adversarial loss for perceptual compression. (2) A diffusion model simultaneously models all the sources’ latents concatenated together as one latent. Inference includes two sub-tasks as well: (1) Total generation allows unconditional generation of all instrumental sources at the same time and (2) Partial generation allows the generation of companion sources given any combinations of instrumental sources (e.g., generate bass and guitar to accompany given piano and drums).
@@ II-A SourceVAE
The SourceVAE aims to compress waveform-domain instrumental sources into a compact latent space, while still ensuring perceptually indistinguishable reconstruction. This is usually achieved by adversarial training with carefully designed discriminators. We borrow this training framework and model architecture from the DAC [10] neural audio codec. DAC is a state-of-the-art waveform-domain neural audio codec trained with both reconstruction and adversarial losses. It can encode, quantize, and decode audio with superior quality. However, we want our latent to be noise-robust and continuous, so we remove the vector quantization module, constrain the intermediate latent size, and add a small KL-divergence loss term as used in vanilla VAEs [33]. The KL-divergence loss is to ensure a noise-robust latent space. For the encoder and decoder, we use DAC’s default 24kHz model but with a new latent size of $C=80$. Given an instrumental source $s\in\mathbb{R}^{N}$ with $N$ samples, the encoder encodes the waveform to a posterior $\Psi_{enc}(s)=\mathcal{N}(\cdot|\mu_{z}(s),\Sigma_{z}(s))$, where $\mu_{z}(s)\in\mathbb{R}^{C\times\frac{N}{D}}$ is the posterior mean of the latent and $\Sigma_{s}$ is the corresponding posterior covariance. $D$ is the time-domain downsampling factor of the encoder, which is $320$ in DAC. Then any $z\sim\mathcal{N}(\cdot|\mu_{z}(s),\Sigma_{z}(s))$, processed by the decoder $\psi_{\text{dec}}$, should reconstruct $s$ with good quality. To extract latent features, we take the posterior mean $z_{s}=\mu_{z}(s)$. During training, the loss is as shown below:
@@
where $L_{\text{Mel}}$, $L_{\text{feature}}$, $L_{\text{adversarial}}$ are Mel-reconstruction loss, feature matching loss, and adversarial loss, respectively, as in the DAC training framework. Also $\lambda_{1}=15$, $\lambda_{2}=2$, $\lambda_{3}=1$ according to DAC. $L_{\text{KL}}$ is the KL-Divergence between the VAE encoded posterior and $\mathcal{N}(0,I)$ and we set $\lambda_{4}=10$. The implementation details are available in our source code.
@@ II-C1 Total Generation
The total generation inference is straightforward. Starting from randomly sampled white noise $Z(T)\sim\mathcal{N}(\cdot|0,\sigma_{max}^{2}I)$, the diffusion sampling process gradually transforms $Z(T)$ to $Z(0)$. Then instrumental source latent $z_{1},z_{2},...,z_{K}$ are extracted from $Z(0)$, and are further decoded independently by the SourceVAE decoder to get the generated source waveforms $\{s_{i}\in\mathbb{R}^{N}|s_{i}=\psi_{dec}(z_{i}),i\in[1,2,...,K]\}$. These generated sources could then be added to form a mixture of music pieces $x=\sum_{k=1}^{K}s_{i}$.
@@ II-C2 Partial Generation
Partial generation is the task of generating complementary sources given some existing ones to condition on. Assume a subset of instruments are given by the indices $I\subset\{1,2,...K\}$, and the corresponding given sources are denoted by $S_{I}=\{s_{i}\}_{i\in I}$. Then the complementary sources to generate are indexed by $\bar{I}=\{1,...,K\}\backslash I$, which means the source to generate are $S_{\bar{I}}=\{s_{i}\}_{i\in\bar{I}}$. Since our diffusion model works in the latent domain, we first use SourceVAE to encode each source in $S_{I}$ to latent $Z_{I}=\{z_{i}|z_{i}=\mu_{z}(s_{i}),i\in I\}$. The task is to generate $Z_{\bar{I}}$ conditioned on $Z_{I}$, so we need the conditional score $\nabla_{{Z}_{\bar{I}}(t)}\log p(Z_{\bar{I}}(t)|Z_{I}(t))$ for sampling. Following diffusion-based imputation [34] and MSDM [29], we could estimate the condition score by:
@@
$d{Z_{\bar{I}}}(t)=\sigma(t)S^{\theta}([Z_{\bar{I}}(t),\hat{Z}_{I}(t)],\sigma(t))dt$ (6)
With ${Z_{\bar{I}}}(0)$ sampled, the partially generated sources $S_{\bar{I}}$ can be decoded by the SourceVAE decoder: $S_{\bar{I}}=\{s_{i}\in\mathbb{R}^{N}|s_{i}=\psi_{dec}(z_{i}),z_{i}\in Z_{\bar{I}}\}$. Then all the conditional sources and partially generated sources are added to form the final music piece $x$.
@@ III-A Dataset
We use the same dataset as MSDM [29], namely the slakh2100 music dataset [36]. slakh2100 is a MIDI synthesized music dataset with 145 hours of music, containing both the mixed music and the individual tracks labeled with instrument class. Same as MSDM, we use $K=4$ main tracks which are piano, drums, bass, and guitar for multi-source modeling. For fair comparison, we use the identical sampling rate of 22,050Hz.
@@ III-B SourceVAE
The SourceVAE mentioned in Sec. II-A is a 1D-CNN-based encoder-decoder architecture coupled with a DAC loss and a KL-divergence loss. The encoder and decoder all follow the final setup of the DAC architecture for 24kHz. The intermediate latent dimension for SourceVAE is set to be $C=80$ and the encoder has a downsampling rate of $D=320$ for the temporal dimension. For training, we use a batch size of 28 and train on one-second-long single-instrumental segments for 100k steps. All other SourceVAE training configurations are the same as the DAC paper with code available at https://github.com/descriptinc/descript-audio-codec.
@@ III-C Latent Diffusion and Unet Architecture
To accommodate the 1D-Unet architecture used in MSDM, we concatenate the K source latent channel-wise, so the new channel dimension becomes $KC$, and the input to the 1D-Unet is $Z^{\prime}(t)\in\mathbb{R}^{KC\times\frac{N}{D}}$, which is a reshaped version of $Z(t)$. $KC$ is treated as the channel dimension and $\frac{N}{D}$ is treated as the temporal dimension for the Unet. In our setup, $K=4,C=80,D=320$. We train our diffusion model on segments with $N=327672$ samples (about 15 seconds). For the 1D-Unet architecture, similar to MSDM, we adapt the architecture used in Moûsai [22] but with some modifications. We set the input channel dimension to be $KC=320$ and then we experiment on two different configurations. We call one model MSLDM and one larger model MSLDM-Large. The MSLDM contains 6 nested U-Net blocks with increasing channels [1024, 2048, 4096, 4096, 4096, 4096]. The downsampling factor for the blocks is [1,1,2,1,1,2]. The self-attention blocks are used for all the blocks except the first one. 12 Attention heads are used, and each head is 64 dimensional. For MSDLM-Large, the Unet contains 8 layers where the corresponding output channel dimensions are [1024, 2048, 4096, 4096, 4096, 4096, 4096, 4096]. The corresponding downsampling factors are [1,1,2,1,1,2,2,2]. All blocks contain self-attention blocks (except the first one) and each attention block contains 12 attention heads that are each 128 dimensional. The Unet and diffusion code setup are adapted from audio-diffusion-pytorch. Similar to MSDM, we train the diffusion model with a batch size 16 and a learning rate of 2e-5 for 400k steps.
All the model training is performed on a single RTX A6000 GPU with 48GB of VRAM. Further details can be found in our code.
@@ IV Evaluation Metrics and Results
We evaluate two tasks: (1) Total generation and (2) Partial generation as mentioned in Sec. II-C. For both tasks, we evaluate the FAD score [37] and human subjective score with a listening test. We compare our performance against 3 baselines.
@@ IV-A Baseline Models
The first baseline is the MSDM [29] model. To show the effectiveness of modeling sources instead of mixtures, we design another baseline called MixLDM, where the latent diffusion model directly models the latent of music mixture, instead of instrumental sources. This is a more common practice in diffusion-based audio/music generation, where the model directly models the music mixture. For training this model, we first train a MixtureVAE which is the same as SourceVAE except that it is trained on Mixture music. The latent size $C=320$, so that the diffusion model’s input is of the same dimension as MSLDM. To claim that our model generates sources that are mutually coherent (i.e. in harmony with each other), we design one baseline called ISLDM (Independent Source Latent Diffusion Model), where we train four independent diffusion models on four instruments’ latents, respectively, so each model can generate one single instrument. For each single source model, the input channel size becomes $C=80$, and all the other training parameters are the same as MSLDM. Since all the single source models are independent of each other, they cannot generate mutually coherent sources. This is set as a baseline to validate our model’s abilities to generate mutually coherent sources. All models’ parameters and inference time to generate one 12-second mixture on a single RTX A6000 GPU are listed in Table I.
@@
MSLDM-Ours 364 5.44
MSLDM-Large-Ours 1654 7.44
@@
TABLE II: sub-FAD for Partial Generation. The sub-FAD (lower is better) is reported for any source combinations (B: Bass, D: Drums, G: Guitar, P: Piano), where BD means conditioned on piano and guitar, the task is to generate Bass and Drums.
@@
IV-B Fréchet Audio Distance
(FAD) We use the Fréchet Audio Distance (FAD) [37] with VGGish feature [38] as the objective metric to evaluate total generation and partial generation. We use 200 chunks of music segment for the FAD calculation, where each segment is about 12 seconds, as in MSDM [29]. We randomly sample 200 12-second long music segments from the slakh2100 test set as the reference set.